OpenStudy (rock_mit182):

Simplify your answer as much as possible. Muliply (3-5*sqroot(2)) (2*sqroot(10) + 4)

OpenStudy (rock_mit182):

\[(3-5\sqrt{2})(2\sqrt{10} + 4)\]

OpenStudy (rock_mit182):

\[6\sqrt{10}-20(\sqrt{5}+\sqrt{2}) + 12\] right ?

OpenStudy (rock_mit182):

Dude Im jus checking my answer

OpenStudy (rock_mit182):

am I right

OpenStudy (rock_mit182):

Or it could be simplify even more ?

OpenStudy (jdoe0001):

\(\bf (3-5\sqrt{2})(2\sqrt{10}+4)\implies 3\cdot 2\sqrt{10}+3\cdot 4-5\sqrt{2}\cdot 2\sqrt{10}-5\sqrt{2}\cdot 4 \\ \quad \\ 6\sqrt{10}+12-10\sqrt{2\cdot 10}-20\sqrt{2}\implies 6\sqrt{10}+12-10\sqrt{20}-20\sqrt{2} \\ \quad \\ {\color{brown}{ 20\to 2\cdot 2\cdot 5\to 2^2\cdot 5}} \\ \quad \\ 6\sqrt{10}+12-10\sqrt{2^2\cdot 5}-20\sqrt{2}\implies 6\sqrt{10}+12-20\sqrt{ 5}-20\sqrt{2}\)

OpenStudy (rock_mit182):

well i guess is right the only difference is the common factor

OpenStudy (jdoe0001):

yeap

OpenStudy (rock_mit182):

Thanks again!

OpenStudy (jdoe0001):

yw